# Recommender Systems for Good (RS4Good): Survey of Use Cases and a Call to Action for Research that Matters

### v1

@@ Abstract.
In the area of recommender systems, the vast majority of research efforts is spent on developing increasingly sophisticated recommendation models, also using increasingly more computational resources. Unfortunately, most of these research efforts target a very small set of application domains, mostly e-commerce and media recommendation. Furthermore, many of these models are never evaluated with users, let alone put into practice. The scientific, economic and societal value of much of these efforts by scholars therefore remains largely unclear. To achieve a stronger positive impact resulting from these efforts, we posit that we as a research community should more often address use cases where recommender systems contribute to societal good (RS4Good). In this opinion piece, we first discuss a number of examples where the use of recommender systems for problems of societal concern has been successfully explored in the literature. We then proceed by outlining a paradigmatic shift that is needed to conduct successful RS4Good research, where the key ingredients are interdisciplinary collaborations and longitudinal evaluation approaches with humans in the loop.
@@
copyright: acmlicensedjournalyear: 2024journal: TORSjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Information systems Recommender systems
@@ 1. Introduction
The question arises why we as a recommender systems research community continue to put so much effort into such a small set of application domains and delimit ourselves to only a few research methodologies. Actually, the literature contains a multitude of examples where recommender systems are applied to a variety of application domains beyond movies and e-commerce. Moreover, the value that is created in several of these settings goes beyond the creation of economic value for the platform provider as well as beyond individual hedonic or utilitarian needs of consumers. We may, for instance, think of recommender systems that guide users to a healthier life style, recommender systems that stimulate energy-saving behavior, or systems that help users achieve self-actualization goals.
Given that these application areas are largely underexplored in the literature, we call the research community to re-focus our combined efforts to areas where recommender systems can contribute to the achievement of societal goals. Following ideas of recent initiatives like ‘AI4Good’ [^2] https://ai4good.org/ or ‘IR4Good’ [^3] https://www.ecir2024.org/2023/07/24/call-for-ir-for-good-papers/, we envision the development of an ‘RS4Good’ movement in the recommender systems community. As a result, we hope that the future research efforts in our community will be better aligned with the pressing challenges that we face as a society. However, as we will lay out in more depth in the following sections, to achieve these goals it will not be sufficient to only use different datasets, but a paradigmatic shift is required, instead. It involves the following components: (i) a focus on problems of societal relevance, (ii) a multi-disciplinary research approach involving humans in the loop, and, (iii) the consideration of the longitudinal effects of recommender systems.
@@ 2.1. Defining RS4Good: From Avoiding Harm to Doing Good
Such avoidance or mitigation strategies are different from our notion of RS4Good that we are advocating in this paper. Rather than to avoid problems that may for example be caused by the underlying business goals or by the unintended reinforcement of biases in the underlying data, social benefit is part of the ‘genes’ and main purpose of an RS4Good recommender system. Certainly, this does not contradict the assumption that recommender systems are designed to create economic value for providers.
Let us think of an alternative social media platform that tries to proactively balance the coverage of content from different political viewpoints, thereby implementing a corporate mission to reduce radicalization and polarization in society. However, such a societal mission does not necessarily mean that economic goals may not be pursued in parallel. The provision of balanced and diversified content may in fact be a key factor of positioning the platform on the market, similar to independent publishers or traditional newspapers in media environments.
@@ 2.2. Selected Use Cases of RS4Good Research
In their recent work, Felfernig et al. (Felfernig et al., 2023) discussed the potential role of recommender systems to support the United Nation’s 17 Sustainability Development Goals (SDGs). These SDGs comprise both areas that are well aligned with typical goals of recommender systems (e.g., ‘economic growth’) as well as areas which, as of today, are barely related to existing recommender systems research (e.g., ‘life below water’). In our following discussion of the RS4Good research landscape, we will therefore not follow the UN’s SDG organization, but rather structure our work along application domains where the use of recommendation systems has been already explored at least to a certain extent. We emphasize that our discussion and categorization is not meant to be exhaustive. Instead, it shall serve as an illustration and motivation for the many possible directions that our community should explore in more depth in the future.
We note that our discussion of existing works is mainly organized in terms of application domains. We however recall that whether a recommender systems is considered to be ‘for good’ does not depend on the application domain. A recommendation system in the e-commerce domain, for example, can be guided purely by economic objectives. It may however also have elements of an RS4Good system, e.g., when the recommender is designed to prioritize sustainable product alternatives. As mentioned in the previous section, it depends on the underlying goals that are implemented in the recommender system. In the following, we review a number of RS4Good systems in different application contexts. Figure 1 illustrates the main topic areas. [^4] We deliberately leave out the general goal of economic growth in this list while acknowledging that the growth of individual businesses and the economy may generally contribute to positive societal developments.
@@ 2.2.1. Health
The health domain is a prime example where recommender systems can be applied for societal good in various ways. A survey of recommender systems in healthcare can be found in (Tran et al., 2021). In this survey, the authors identify the following main application areas.
@@
A number of works in the literature cover additional topics that go beyond this classification, e.g., the recommendation of messages within a smoking cessation app (Hors-Fraile et al., 2016). Such works fall into the more general category of behavior change support systems (Oinas-Kukkonen, 2010). Some of the works in the literature are also targeted at certain groups of people or patients, for instance, elderly (Espín et al., 2016; Herpich et al., 2017; Hammer et al., 2015) or people with autism (Mauro et al., 2022), dementia (Steinert et al., 2022), or social anxiety (Ameko et al., 2020). Today, many of the more recent health-related recommendation approaches are implemented as (smartphone) apps. This leads to additional research questions beyond those related to the core recommendation task. The credibility of a health-related app is for example, important for the adoption of the recommendation, as analyzed in (Oyibo et al., 2018). Also there are often ethical aspects to consider, in particular in the context of mental health apps (Valentine et al., 2023; Rohani et al., 2020). Of course, it is often highly important to limit the risks of inadequate recommendations in such a critical domain. This can for example be achieved by providing adequate explanations alongside recommendations, and by leaving the final decision of the chosen health intervention to the medical expert (de Lima et al., 2021).
@@ 2.2.2. Educational Recommender Systems and e-Learning
Another quite active and established application domain for RS4Good is the use of recommendation technology to support human/student learning processes, e.g., in the form of course recommendations or learning paths suggestions in e-learning environments. Works in this area fall into the broader category of ‘AI in education’, which is of UNESCO’s focus topics [^6] https://www.unesco.org/en/digital-education/artificial-intelligence. Also, e-learning recommendations represent one form of Technology-enhanced Learning (TEL) (Duval et al., 2017).
Several types of items can be recommended in e-learning recommender systems, including courses and course paths, college majors, papers, web pages, lessons and various other learning resources. Various types of traditional recommendation algorithms—both content-based, collaborative filtering and hybrid ones—have been explored in the literature (Klašnja-Milićević et al., 2015). In some approaches, the recommendation of learning activities (and materials) is based on predicting the student’s performance on a task that she or he has not solved yet. Given these predictions, the set of recommendations is then determined based on the e-learning system’s specific goals. For example, the system could recommend materials that the student assumedly will like and most probably will be able to master. Following the exact opposite strategy, the system may however also recommend materials where it assumes gaps in the student’s knowledge.
An in-depth discussion of the challenges of building e-learning recommender systems can be found in (Klašnja-Milićević et al., 2015). A recommender system not only has to understand the learning goals and the prior knowledge of the student, it has also to take into account personal characteristics, such the student’s learning strategy. Moreover, recommendations must be made at the right point in time and within an appropriate interactive tool environment. Generally, technology-enhanced learning processes and educational recommender systems are a multi-faceted topic, and research in this direction usually requires an interdisciplinary approach (Balacheff et al., 2009). Furthermore, the evaluation of e-learning recommender systems remains difficult. As mentioned above, depending on the system’s goal and strategy, one or the other type of learning content may be appropriate and a good recommendation for a given student. Furthermore, the true success of the system, i.e., if the student’s knowledge or learning process, was improved, can usually only be evaluated from a longitudinal perspective.
@@ 2.2.3. Individual Development, Wellbeing and Safety
Recreational sport is often considered to be an important contributor to wellbeing. The use of recommendation technology and machine learning techniques for amateur marathon runners was for example proposed in Smyth et al. (Smyth et al., 2022). The primary recommendation tasks in their approach range from the suggestion of personalized training plans to pacing strategies and the prediction of realistic finish-times. Secondary use cases for runners include the selection of running events, running routes, training partners, gear and equipment as well as advice regarding nutrition and recovery. In (Smyth et al., 2022), a number of these uses cases for recommendation technology are discussed in more depth. The evaluation of the proposed approaches was however largely limited to offline experiments so far.
Going beyond the realm of sports and health, Knijnenburg et al. (Knijnenburg et al., 2016) propose to develop a new class of recommender systems for self-actualization. Instead of just recommending the “best” items to users, such systems are designed to “support users in developing, exploring, and understanding their own unique tastes and preferences.” As a consequence, future recommender systems for self-actualization should help users break out of filter bubbles and rather support instead of replace them in their decision-making tasks. A possible experiment to evaluate such a recommender system was later outlined in (Wilkinson et al., 2017). The results from a user study in the news recommendation domain are discussed in (Sullivan et al., 2019). In this study, it was found that the reading intention of the participants are influenced by their stated self-actualization goals, which could either be Broaden Horizons or Discover the Unexplored.
@@ 2.2.4. Societal and Political Development
Also in the realm of social media, personalization and recommender systems are often assumed to be the cause of echo chambers, filter bubbles and polarization (Donkers and Ziegler, 2021; Stray, 2022). Donkers et al. (Donkers and Ziegler, 2021) investigate the phenomenon of ‘dual’ echo chambers with an agent-based modeling and simulation approach. Their study reveals that counteracting the two types of echo chambers requires different proactive diversification and depolarization strategy. In a similar vein, Stray (Stray, 2022) studies how to design recommender systems for depolarization. Diversification of the presented content is considered as one possible strategy that can work against polarization in some context, but it may even make things worse. According to (Stray, 2022), recommenders should generally be designed to deprioritize content that is found to be polarizing as well as polarization measures should be incorporated into systems and continuously monitored.
@@
Finally, we mention two examples of works that may lead to a positive and socially desirable economic development in quite different ways. First, Pourashraf and Mobasher (Pourashraf and Mobasher, 2022) target the news recommendation domain and argue that readers’ local news interests are different from their preferences for global news. Thus, by differentiating between local and global news preferences their localized recommendation model increased prediction accuracy with the ultimate goal to revitalize and strengthen local media companies. The second work from Liu et al. (Liu et al., 2019) target the domain of micro-lending, an approach to provide citizens in impoverished countries with access to capital. In their work, the authors propose a fairness-aware model that aims to ensure that different demographic groups have a fair chance of being recommended for a micro-loan.

### v2

@@ Abstract.
In the area of recommender systems, the vast majority of research efforts is spent on developing increasingly sophisticated recommendation models, also using increasingly more computational resources. Unfortunately, most of these research efforts target a very small set of application domains, mostly e-commerce and media recommendation. Furthermore, many of these models are never evaluated with users, let alone put into practice. The scientific, economic and societal value of much of these efforts by scholars therefore remains largely unclear. To achieve a stronger positive impact resulting from these efforts, we posit that we as a research community should more often address use cases where recommender systems contribute to societal good (RS4Good). In this opinion piece, we first discuss a number of examples where the use of recommender systems for problems of societal concern has been successfully explored in the literature. We then proceed by outlining a paradigmatic shift that is needed to conduct successful RS4Good research, where the key ingredients are interdisciplinary collaborations and longitudinal evaluation approaches with humans in the loop. [^1] Paper accepted for publication in ACM Transactions on Recommender Systems
@@
copyright: acmlicensedjournalyear: 2025doi: 10.1145/3746648journal: TORSjournalvolume: 1journalnumber: 1article: 1publicationmonth: 1ccs: Information systems Recommender systems
@@ 1. Introduction
For us as a recommender systems research community, this raises the following question: why do we continue to focus so narrowly on a few application domains and methodologies? Actually, the literature contains a multitude of examples where recommender systems are applied to a variety of application domains beyond movies and e-commerce. Moreover, the value that is created in several of these settings goes beyond the creation of economic value for the platform provider as well as beyond individual hedonic or utilitarian needs of consumers. We may, for instance, think of recommender systems that guide users to a healthier life style, recommender systems that stimulate energy-saving behavior, or systems that help users achieve self-actualization goals.
Given that these application areas are largely underexplored in the literature, we call the research community to re-focus our combined efforts to areas where recommender systems can contribute to the achievement of societal goals. Following ideas of recent initiatives like ‘AI4Good’ [^3] https://ai4good.org/ or ‘IR4Good’ [^4] https://www.ecir2024.org/2023/07/24/call-for-ir-for-good-papers/, we envision the development of an ‘RS4Good’ movement in the recommender systems community. As a result, we hope that the future research efforts in this area will be better aligned with the pressing challenges that we face as a society. However, as we will lay out in more depth in the following sections, to achieve these goals it will not be sufficient to only use different datasets, but a paradigmatic shift is required, instead. It involves the following components: (i) a focus on problems of societal relevance, (ii) a multi-disciplinary research approach involving humans in the loop, and, (iii) the consideration of the longitudinal effects of recommender systems.
@@ 2.1. Defining RS4Good: From Avoiding Harm to Doing Good
Such avoidance or mitigation strategies are different from our notion of RS4Good that we are advocating in this paper. Rather than to avoid problems that may for example be caused by the underlying business goals or by the unintended reinforcement of biases in the underlying data, social benefit is part of the ‘genes’ and main purpose of an RS4Good recommender system. Certainly, this does not contradict the assumption that recommender systems are designed to create economic value for providers. Let us think of an alternative social media platform that tries to proactively balance the coverage of content from different political viewpoints, thereby implementing a corporate mission to reduce radicalization and polarization in society. However, such a societal mission does not necessarily mean that economic goals may not be pursued in parallel. The provision of balanced and diversified content may in fact be a key factor of positioning the platform on the market, similar to independent publishers or traditional newspapers in media environments.
@@ 2.2. Selected Use Cases of RS4Good Research
In their recent work, Felfernig et al. (Felfernig et al., 2023) discussed the potential role of recommender systems to support the United Nation’s 17 Sustainability Development Goals (SDGs). These SDGs comprise both areas that are well aligned with typical goals of recommender systems (e.g., ‘economic growth’) as well as areas which, as of today, are barely related to existing recommender systems research (e.g., ‘life below water’). In our following discussion of the RS4Good research landscape, we will therefore not follow the UN’s SDG organization, but rather structure our work along application domains where the use of recommendation systems has been already explored at least to a certain extent. We emphasize that our discussion and categorization is not meant to be exhaustive. Instead, it shall serve as an illustration and motivation for the many possible directions that the research community should explore in more depth in the future.
We note that our discussion of existing works is mainly organized in terms of application domains for presentation purposes. We however recall that whether a recommender systems is considered to be ‘for good’ does not necessarily depend on the application domain. A recommendation system in the e-commerce domain, for example, can be guided purely by economic objectives (Zanker et al., 2006). It may however also have elements of an RS4Good system, e.g., when the recommender is designed to prioritize sustainable product alternatives. As mentioned in the previous section, it depends on the underlying goals that are implemented in the recommender system. In the following, we review a number of RS4Good systems in different application contexts. Figure 1 illustrates the main topic areas. [^5] We deliberately leave out the general goal of economic growth in this list while acknowledging that the growth of individual businesses and the economy may generally contribute to positive societal developments. Other categorizations of existing works are possible as well, e.g., by distinguishing between individual and societal goals. We discuss questions of goal identification and goal setting later in Section 3.2.
@@
We re-iterate that our discussions of individual research works in each application area are not intended to represent a comprehensive review of the partially quite rich literature, e.g, in the educational domain, but serve as motivational examples. An overview of selected in-depth surveys in relevant domains can be found in Table 1.
Table 1. Selected survey works in different domains.
Domain References
Health and Healthcare (De Croon et al., 2021), (Etemadi et al., 2023), (Pincay et al., 2019), (Tran et al., 2021)
Education and e-learning (Bodily and Verbert, 2017), (da Silva et al., 2023), (Klašnja-Milićević et al., 2015), (Urdaneta-Ponte et al., 2021)
Individual Development & Well-Being (Elsweiler et al., 2022), (Felfernig et al., 2024), (Mahajan and Kaur, 2024), (Tran et al., 2018)
Societal and Political Development (Bauer et al., 2024), (Kamal et al., 2024)
Environment Protection & Sustainability (Felfernig et al., 2023), (Zhou et al., 2024)
@@ 2.2.1. Health
The health domain is a prime example where recommender systems can be applied for societal good in various ways. One notable survey of recommender systems in healthcare can be found in (Tran et al., 2021). In this survey, the authors identify the following main application areas.
@@
A number of works in the literature cover additional topics that go beyond this classification, e.g., the recommendation of messages within a smoking cessation app (Hors-Fraile et al., 2016). Such works fall into the more general category of behavior change support systems (Oinas-Kukkonen, 2010). Some of the works in the literature are also targeted at certain groups of people or patients, for instance, elderly (Espín et al., 2016; Herpich et al., 2017; Hammer et al., 2015) or people with autism (Mauro et al., 2022), dementia (Steinert et al., 2022), or social anxiety (Ameko et al., 2020). Today, many of the more recent health-related recommendation approaches are implemented as (smartphone) apps. This leads to additional research questions beyond those related to the core recommendation task. The credibility of a health-related app is for example, important for the adoption of the recommendation, as analyzed in (Oyibo et al., 2018). Also there are often ethical aspects to consider, in particular in the context of mental health apps (Valentine et al., 2023; Rohani et al., 2020). Of course, it is often highly important to limit the risks of inadequate recommendations in such a critical domain. This can, for instance, be achieved by providing adequate explanations alongside recommendations, and by leaving the final decision of the chosen health intervention to the medical expert (de Lima et al., 2021).
@@ 2.2.2. Educational Recommender Systems and e-Learning
Another quite active and established application domain for RS4Good is the use of recommendation technology to support human/student learning processes, e.g., in the form of course recommendations or learning paths suggestions in e-learning environments. Works in this area fall into the broader category of ‘AI in education’, which is one of UNESCO’s focus topics [^7] https://www.unesco.org/en/digital-education/artificial-intelligence. Also, e-learning recommendations represent one form of Technology-enhanced Learning (TEL) (Duval et al., 2017).
Several types of items can be recommended in e-learning recommender systems, including courses and course paths, college majors, papers, web pages, lessons and various other learning resources. Various types of traditional recommendation algorithms—both content-based, collaborative filtering and hybrid ones—have been explored in the literature (Klašnja-Milićević et al., 2015). In some approaches, the recommendation of learning activities (and materials) is based on predicting the student’s performance on a task that she or he has not solved yet. Given these predictions, the set of recommendations is then determined based on the e-learning system’s specific goals. For example, the system could recommend materials that the student assumedly will like and will most probably be able to master. Following the exact opposite strategy, the system may however also recommend materials where it assumes gaps in the student’s knowledge. In this context, it is important to note that recommending only courses and materials for which the system assumes that the student will receive a good grade may actually be counterproductive, e.g., when the system starts recommending only easy courses. Instead, one could argue that recommending more challenging courses—or courses that lead to a well-rounded overall educational profile—will contribute better to a student’s growth and development.
An in-depth discussion of the challenges of building an e-learning recommender systems can be found in (Klašnja-Milićević et al., 2015). A recommender system not only has to understand the learning goals and the prior knowledge of the student, it has also to take into account personal characteristics, such as the student’s learning strategy. Moreover, recommendations must be made at the right point in time and within an appropriate interactive tool environment. Generally, technology-enhanced learning processes and educational recommender systems are a multi-faceted topic, and research in this direction usually requires an interdisciplinary approach (Balacheff et al., 2009). Furthermore, the evaluation of e-learning recommender systems remains difficult. As mentioned above, depending on the system’s goal and strategy, one or the other type of learning content may constitute appropriate recommendations for given students and their assumed educational goals. Furthermore, the true success of the system, i.e., if the student’s knowledge or learning process, was improved, can usually only be evaluated from a longitudinal perspective.
@@ 2.2.3. Individual Development, Wellbeing and Safety
In this context, we acknowledge that some of the works discussed here, i.e., those that target individual wellbeing, could also be seen to fall under the category of health recommender systems discussed above. In the individual wellbeing category, we consider works that do not focus primarily on the clinical aspects of health or contribute to health as a societal goal. Instead, the works discussed here are often about behavior change, prevention, or individual goal-setting. Consider, for example, the work by Radha et al. (2016). In this research, the authors investigate a patient-centered recommender system that provides personalized lifestyle modification advice for blood pressure management, considering also the individual patient’s ability level. Similarly, the user’s ability level and task difficulties were also taken into account in a nutrition recommender systems in (Schäfer and Willemsen, 2019) and in a behavior-change recommender system in (Torkamaan and Ziegler, 2022).
Recreational sport is often considered to be another important contributor to wellbeing. The use of recommendation technology and machine learning techniques for amateur marathon runners was for example proposed in Smyth et al. (Smyth et al., 2022). The primary recommendation tasks in their approach range from the suggestion of personalized training plans to pacing strategies and the prediction of realistic finish-times. Secondary use cases for runners include the selection of running events, running routes, training partners, gear and equipment as well as advice regarding nutrition and recovery. In (Smyth et al., 2022), a number of these uses cases for recommendation technology are discussed in more depth. The evaluation of the proposed approaches was however largely limited to offline experiments so far.
Going beyond the realm of sports and health, Knijnenburg et al. (Knijnenburg et al., 2016) propose to develop a new class of recommender systems for self-actualization. Instead of just recommending the “best” items to users, such systems are designed to “support users in developing, exploring, and understanding their own unique tastes and preferences.” As a consequence, future recommender systems for self-actualization should help users break out of filter bubbles and rather support users instead of replacing them in their decision-making tasks. A possible experiment to evaluate such a recommender system was later outlined in (Wilkinson et al., 2017). The results from a user study in the news recommendation domain are discussed in (Sullivan et al., 2019). In this study, it was found that the reading intention of participants is influenced by their stated self-actualization goals, which could either be Broaden Horizons or Discover the Unexplored. In particular in this area of self-actualization and personal development, it generally seems highly important to properly understand the user’s goals, and to actively involve them in the recommendation process instead of just interpreting their observed behavior (Ekstrand and Willemsen, 2016; Liang et al., 2023).
@@ 2.2.4. Societal and Political Development
Also in the realm of social media, personalization and recommender systems are often assumed to be the cause of echo chambers, filter bubbles and polarization (Donkers and Ziegler, 2021; Stray, 2022). Donkers et al. (Donkers and Ziegler, 2021) investigate the phenomenon of ‘dual’ echo chambers with an agent-based modeling and simulation approach. Their study reveals that counteracting the two types of echo chambers requires different proactive diversification and depolarization strategy. In a similar vein, Stray (Stray, 2022) studies how to design recommender systems for depolarization. Diversification of the presented content is considered as one possible strategy that can work against polarization in some context, but it may even make things worse. According to (Stray, 2022), recommenders should generally be designed to deprioritize content that is found to be polarizing, and polarization measures should be incorporated into systems and continuously monitored. In that context, Ovadya and Thorburn (2023) propose what they call bridging systems: systems that promote mutual understanding and trust and which thereby help reducing polarization. Among other examples, the authors discuss how recommender systems on social media could be designed to implement a bridging functionality. Instead of designing the system to maximize engagement, which may be achieved by recommending controversial content, a bridging recommender should for example emphasize content which is endorsed by people who in the past had diverse viewpoints.
@@
Finally, we mention two examples of works that may lead to a positive and socially desirable economic development in quite different ways. First, Pourashraf and Mobasher (Pourashraf and Mobasher, 2022) target the news recommendation domain and argue that readers’ local news interests are different from their preferences for global news. Thus, by differentiating between local and global news preferences their localized recommendation model increased prediction accuracy with the ultimate goal to revitalize and strengthen local media companies. The second work from Liu et al. (Liu et al., 2019) target the domain of micro-lending, an approach to provide citizens in impoverished countries with access to capital. In their work, the authors propose a fairness-aware model that aims at ensuring that different demographic groups have a fair chance of being recommended for a micro-loan.
@@ 2.2.5. Environment Protection and Sustainability
2.3. RS4Good and General Decision Support Systems
At this point, it is important to iterate what we consider central features of a recommender system and in which ways such systems are different from the class of more general decision-support and advice-giving systems like those just mentioned for the agricultural domain. First, in a recommender system, there commonly is the underlying goal to help users in situations of information overload, where users are confronted with a relatively rich set of options to choose from. Second, we assume that recommender systems are designed to personalize their suggestions based on the needs, preferences and other characteristics of individual users. Third, users are assumed to chose individual items from a given set of fixed or preconfigured alternatives, such as picking a movie to watch.
Given this characterisation of the key features of a personalized recommender system, a crop yield prediction system, as for example presented in (Pande et al., 2021), would not fall under our definition even though called recommender system by the authors. Likewise, there are various other works, e.g., in the health domain, which are sometimes referred to as recommender systems, but are not covered by our definition. The famous MYCIN system (Shortliffe and Buchanan, 1984) from the 1980s, for example, was an expert system implementing hundreds of explicit rules. These rules allowed the system not only to analyze a patient’s symptoms, but they could also recommend antibiotics. Nonetheless, we would not consider MYCIN to be a recommender system, as it does not consider the preferences of the physician or the patient. Let us note, however, that the boundary between personalized recommender systems and the more general class of advice-giving systems is sometimes blurred. [^9] Conversational and knowledge-based recommender systems, for example, typically do not rely on past user preferences. In bundle recommendation scenarios, a recommendation refers to a set of multiple and dynamically combined items. And in future scenarios based on Generative AI, the items may dynamically be created or tailored according to user preferences.
Certainly, systems like the mentioned agricultural advice-giving solution and medical decision-support systems are important tools that can help promote societal good. Nonetheless, with this essay we primarily aim to address researchers who focus on personalized recommender systems. [^10] Non-personalized decision-support systems may be covered by more general ‘AI4Good’ initiatives. We argue that there are plenty of opportunities for such systems also in domains like healthcare and we aim to encourage recommender systems researchers to more often explore the use of personalized solutions in such domains.